# Towards Vision-Language Geo-Foundation Model: A Survey

### v1

@@ 2.4 Related Research Domains
$\bullet$ Vision geo-foundation models. VGFMs are also called remote sensing pre-trained models. The primary difference between these models and VLGFM is that their input is purely visual, without any modeling of textual information. RVSA [61] trained a conventional vision transformer with 100 million parameters on RGB RSIs and developed a novel rotational resizing window attention mechanism to fine-tune the model for downstream tasks. Unlike MAE-based methods that infer the entire image from a few visible patches, RingMo [62] implemented a masked image modelling strategy in their pre-trained model, considering all image patches, whether masked or unmasked. The success of these preliminary studies demonstrates the significant potential of pre-trained models in remote sensing applications. Recently, foundational visual models for spectral [63], multimodal [15], and synthetic aperture radar (SAR) [64] images have been proposed, attracting significant attention in the VGFM field.
$\bullet$ Remote sensing LLM-powered agent. The concept of an LLM-powered agent emphasizes leveraging the capabilities of LLMs to decompose tasks into multiple steps and then accomplish specific tasks by invoking appropriate tools. The most well-known agent framework currently is LangChain [^1] https://www.langchain.com, which allows developers to integrate large language models with external computation and data sources [65, 66, 67, 68]. In the field of remote sensing interpretation, an LLM-powered agent [69] would utilize pre-prepared toolkits, including object detection and classification algorithms, to complete visual-language multimodal tasks. Recently, RS-Agent [70] treats conversational VLGFMs, such as GeoChat and LHRS-Bots, as tools, utilizing them to achieve RSI captioning functionality.
@@ 3 Methods: A survey
TABLE III: Training Data of VLGFMs. Abbreviations IC, VQA, IS, Det., Seg., and VG in the table stand for image captioning, visual question answering, image scene classification, object detection, segmentation and visual grounding, respectively (If more than 1 models use the dataset to evaluate a task, it is highlighted in blue. If more than 2 models use it, it is highlighted in red).
@@
[22] UCM-Captions LEVIR, HRSC, VisDrone,
@@
[25] RSVQA-LR SAMRS
@@
[28] UCM-Captions, RSVQA-LR, DIOR-RSVG, DIOR-Conversa
@@
[29] UCM-Captions, RSVQA-LR, UCM, WHU-RS19, HRRSD, RSOD, UCAS-AOD, VisDrone,
@@
[30] UCM-Caption RSVQA-HR METER-ML,UCM, DIOR-RSVG LHRS-Instruct
@@
[32] MillionAID, NWPU, MSAR, GID, FBP, BANDON,
@@
TABLE IV: Data Pipeline of VLGFMs.
@@
[22] (the first 2 captions are generated according to the target location, and the remaining 3 captions are generated by
@@
[23] image checking and deduplication. Then, clean the dataset using the CLIP score and the RS image classifier. RS3:
@@
[24] type, and resolution; (2) object details such as quantity, color, shape, size, and position; (3) annotate the overall
@@
[25] K-Means. Objects’ relative sizes are small, normal, or large. Objects’ relative positions are assessed using a 3x3
@@
[26] ensuring a uniform location distribution avoiding duplicates, and filtering out indoor photos with a classifier.
@@
[27] (OSM). Connect images with appropriate OSM semantics using a two-stage tag classification approach with
@@
[28] Following the miniGPT-v2, create an phrase grounding dataset. Multi-task Conversation Instruction: mix the
@@
[29] 1 visual grounding dataset into instructions using prompt templates.
@@
[30] keys and balance semantic information. Using Vicuna-v1.5 to generate image captions by summarising the key
@@
[32] remote sensing images, including the type and resolution of the images, as well as information on the attributes,
@@
[34] remote sensing objects.
@@ 3.1 Data Pipeline
$\bullet$ Data collection from scratch. As shown in Table III, three VLGFMs have built their training datasets from scratch rather than utilizing existing ones. RSGPT [24] prioritizes the quality of training samples over their quantity. Hence, it employs five experts to annotate images, producing high-quality captions manually. As a result, the model achieves satisfactory performance with fine-tuning on only 2,500 remote sensing image-text pairs. GRAFT [26] ingeniously utilizes the geo-tag information of images to link ground-level image captions collected from social media with RSIs. This approach enables the acquisition of many remote sensing image-text pairs without the need for manual annotation of RSI captions. SkyScript [27] utilizes a well-known open-source collaborative project, OpenStreetMap [^2] https://www.openstreetmap.org (OSM), to construct image captions. OSM allows users to contribute and edit geospatial data, such as road networks, trails, cafes, railway stations, etc. It provides a comprehensive and freely accessible world map and detailed user annotations. SkyScript filters these annotations and combines various tags to generate high-quality image captions. In LHRS-Bo [30], the LHRS-Align also acquires tags of geographic features from OSM and generates image captions based on these tags using language-only LLM, Vicuna-v1.5. More details can be found in Table IV.
@@
$\bullet$ Data enhancement using existing datasets. Most works choose to modify existing remote sensing datasets, which is significantly less labour-intensive compared to building datasets from scratch. Consequently, researchers can focus more on improving the model’s structures. RemoteCLIP [22] was the first to propose using text templates to convert the most common object detection annotations in the field of remote sensing interpretation into image captions, generating a large number of remote sensing image-text pairs suitable for training large models. Subsequently, EarthGPT [29] also adopted the same approach. RS5M [23], which currently contains the largest image-text pairs, includes 5 million RSIs. It employs the general multimodal model BLIP2 [71] to generate image captions and then uses CLIP [72] to select the top 5 highest-scoring captions. GeoChat [25] constructs a remote sensing image-text pair dataset based on the segmentation dataset generated using SAM [73] technology. It first uses the target’s color, size, and location to build basic expressions and then inputs these expressions into a language-only LLM to generate question-answer pairs. SkyEyeGPT [28] constructs a multi-task conversation instruction dataset by combining an object detection dataset with a VQA dataset built using the object detection data. This approach supports training for both VQA and visual grounding tasks. For more details, please refer to Table IV.
@@ 3.2 Architecture
VLGFMs can be briefly classified into three categories based on their input-output mechanisms. Contrastive VLGFMs [22, 23, 26, 27] take both text and images as inputs and produce a similarity measure between them. This similarity is crucial for image-text retrieval and zero-shot scene classification applications. Conversational VLGFMs [24, 25, 28, 29, 30, 31, 32] also process text and images as inputs but output textual responses, leveraging the powerful language capabilities of LLMs to support tasks such as captioning and visual question answering. Additionally, dense prediction tasks can be supported by appending task-specific heads to the LLM. On the other hand, generative VLGFMs [37, 74] are conditioned on either text or images, producing images as outputs. These models utilize conditional diffusion processes to generate controlled synthetic RSIs. This section will provide a detailed exposition of the model architectures for these three types of VLGFMs.
@@ 3.2.1 Contrastive VLGFMs
A typical contrastive VLGFM comprises two main components: an image encoder and a text encoder. A schematic of this architecture is depicted in Figure 2. Image encoder uses a convolutional neural network or a vision transformer to encode images into a feature vector. Parallel to the vision encoder, the text encoder (typically a BERT model) processes textual descriptions into another feature vector. The core idea behind contrastive VLGFM is to align the image and text embeddings in a shared multidimensional space. During training, contrastive VLGFM uses a contrastive learning objective, specifically a noise contrastive estimation technique variant. The model is trained to maximize the similarity between positive image-text pairs and minimize the similarity between negative pairs. The large-scale training allows contrastive VLGFM to understand various visual concepts described in natural language robustly. CLIP [72] is one of the pioneering models for vision-language pre-training. It employs the InfoNCE loss to align text and images in the feature space. Benefiting from a dataset comprising 400 million image-text pairs, CLIP has achieved significant success in various multimodal downstream tasks. BLIP [75] is the first multimodal model that simultaneously addresses both understanding and generation of image-text content, which is jointly pre-trained on three vision-language tasks: image-text contrastive learning, image-text matching, and image-conditioned language modelling.
In the domain of earth observation, some outstanding vision-language pre-training models have also emerged. RemoteCLIP [22] is the first true contrastive VLGFM. It utilizes existing remote sensing retrieval, detection, and segmentation datasets to construct image-text sample pairs, supporting model training. GeoRSCLIP is a vision-language pre-training model released alongside the RS5M dataset [23]. The authors did not provide extensive details about it in the paper. GRAFT [26] is a significant attempt to address the issue of insufficient samples in remote sensing image-text data. It utilizes a vast collection of ground-level images from Flickr [^3] https://www.flickr.com as a bridge, pairing RSIs that correspond to the geographic coordinates of these ground images with the titles of the ground images. This approach allows for the construction of a large-scale dataset of remote sensing image-text pairs without the need for additional manual annotation. SkyScript [27] utilizes the ground sample distance (GSD) of OSM data to annotate images appropriately. For instance, house numbers cannot be discerned from RSIs regardless of their resolution, and utility poles are not visible in RSIs with a GSD of 10. However, these contrastive VLGFMs are primarily data-centric efforts, with no significant modifications made to the model structures.
@@ 3.2.2 Conversational VLGFMs
Table I presents the visual encoders commonly used in current VLGFM implementations. Besides the standard CLIP image encoder [72], some studies have explored other variants. For instance, RSGPT and SkyEyeGPT employ the EVA-CLIP encoder [76], which is trained with improved training techniques. Additionally, EarthGPT utilizes two encoders simultaneously and introduces the ConvNext encoder [77], a convolution-based architecture designed to exploit higher resolutions and multi-level features.
$\bullet$ Pre-trained LLM. Training an LLM from scratch entails prohibitive computational costs for most individuals, making it more effective and practical to begin with a pre-trained one. Through pre-training on vast amounts of data collected from various places, the LLM acquires rich world knowledge and demonstrates strong capabilities in generalization and reasoning. Table I summarizes commonly used and publicly available LLMs, all of which fall under the category of causal decoders [60]. The LLAMA series [18] and the Vicuna family [78] represent some of the most notable open-source LLMs that have garnered considerable attention within the academic community. It is important to note that these LLMs were primarily pre-trained on English corpora, which limits their multilingual capabilities, such as for Chinese. Additionally, increasing the parameter size of LLMs also brings additional benefits, akin to increasing input resolution. Specifically, RSGPT [24] found that simply expanding the Vicuna from 7B to 13B parameters can significantly enhance performance across various benchmarks.
@@
For token-level connectors, the output of the encoder is transformed into tokens and concatenated with text tokens before being fed into the LLM. A common and feasible solution is to use a set of learnable query tokens to extract information in a query-based manner, initially implemented with BLIP-2 [71] and subsequently adopted by various works on VLGFMs [24]. This query-based approach compresses visual tokens into a smaller number of representation vectors. In contrast, some methods employ multi-layer perception (MLP) to bridge the modal gap. For instance, the LLaVA series only use linear layers [28, 29, 31] or two-layer MLPs [25, 32] to project visual tokens and align feature dimensions with word embeddings.
@@
For feature-level connectors, it is required to incorporate additional modules to facilitate deep interaction and integration between textual features and visual features. For instance, Flamingo [79] introduces additional cross-attention layers between the frozen transformer layers of the LLM, thereby utilizing external visual cues to enhance language features. Similarly, LHRS-Bot [30] attempts to introduce a set of learnable queries for each level of image features. These queries aim to encapsulate the semantic information of each level of visual features through a series of stacked cross-attention and MLP layers.
@@
$\bullet$ Pre-training. Pre-training primarily aims to align different modalities and learn multimodal world knowledge in the initial training phase. The pre-training phase typically requires large-scale image-text pair data, such as caption data. For a given image, the model is trained to autoregressively predict the image’s caption using a standard cross-entropy loss. A common approach during pre-training is to keep pre-trained modules (such as the visual encoder and LLM) frozen and train the vision-language connector [24]. The idea is to align the visual and linguistic modalities without losing the knowledge acquired during pre-training. Some methods [28, 29, 30] also unfreeze LLM to allow more trainable parameters for alignment. ShareGPT4V [80] found that having high-quality caption data during pre-training and unlocking visual encoding capabilities can promote better alignment. In the existing VLGFM landscape, only H2RSVLM [32] has attempted to unfreeze all modules simultaneously for pre-training.
@@
$\bullet$ Instruction-tuning. Instruction-tuning aims to teach models to understand user commands better and complete desired tasks. Through this approach, LLMs can be generalized to support unseen tasks, enhancing zero-shot performance. This simple yet effective idea has spurred the success of subsequent NLP works, such as InstructGPT [81]. Supervised fine-tuning methods typically require much training data to train task-specific models. Instruction-tuning methods reduce the reliance on large-scale data and can accomplish specific tasks through prompt engineering. Due to the flexible format of instructional data and the variability in templates across different tasks, collecting data samples is often more challenging and costly. Recent studies have shown that the quality of instructional fine-tuning samples is more critical than quantity. [82] discovered that high-quality instruction-tuning data can yield better performance, even if the quantity of samples is not the largest.

### v2

@@ 2.4 Related Research Domains
$\bullet$ Vision geo-foundation models. VGFMs are also called remote sensing pre-trained models. The primary difference between these models and VLGFM is that their input is purely visual, without any modeling of textual information. RVSA RVSA trained a conventional vision transformer with 100 million parameters on RGB RSIs and developed a novel rotational resizing window attention mechanism to fine-tune the model for downstream tasks. Unlike MAE-based methods that infer the entire image from a few visible patches, RingMo ringmo implemented a masked image modelling strategy in their pre-trained model, considering all image patches, whether masked or unmasked. The success of these preliminary studies demonstrates the significant potential of pre-trained models in remote sensing applications. Recently, foundational visual models for spectral spectralgpt, multimodal guo2024skysense, and synthetic aperture radar (SAR) l2024saratrx images have been proposed, attracting significant attention in the VGFM field.
$\bullet$ Remote sensing LLM-powered agent. The concept of an LLM-powered agent emphasizes leveraging the capabilities of LLMs to decompose tasks into multiple steps and then accomplish specific tasks by invoking appropriate tools. The most well-known agent framework currently is LangChain [^1] https://www.langchain.com, which allows developers to integrate large language models with external computation and data sources wu2023visual; yang2023mm; shen2024hugginggpt; liu2023interngpt. In the field of remote sensing interpretation, an LLM-powered agent guo2024remoteagent would utilize pre-prepared toolkits, including object detection and classification algorithms, to complete visual-language multimodal tasks. Recently, RS-Agent xu2024rsagent treats conversational VLGFMs, such as GeoChat and LHRS-Bots, as tools, utilizing them to achieve RSI captioning functionality.
@@ 3 Methods: A survey
Table 3: Training Data of VLGFMs. Abbreviations IC, VQA, IS, Det., Seg., and VG in the table stand for image captioning, visual question answering, image scene classification, object detection, segmentation and visual grounding, respectively (If more than 1 models use the dataset to evaluate a task, it is highlighted in blue. If more than 2 models use it, it is highlighted in red).
@@
liu2024remoteclip UCM-Captions LEVIR, HRSC, VisDrone,
@@
kuckreja2023geochat RSVQA-LR SAMRS
@@
zhan2024skyeyegpt UCM-Captions, RSVQA-LR, DIOR-RSVG, DIOR-Conversa
@@
zhang2024earthgpt UCM-Captions, RSVQA-LR, UCM, WHU-RS19, HRRSD, RSOD, UCAS-AOD, VisDrone,
@@
muhtar2024lhrsbot UCM-Caption RSVQA-HR METER-ML,UCM, DIOR-RSVG LHRS-Instruct
@@
pang2024h2rsvlm MillionAID, NWPU, MSAR, GID, FBP, BANDON,
@@
RSCoVLM VHM-Instruct, VHM-Instruct, VHM-Instruct, DOTA refGeo, LLaVA-OneVision
li2025rscovlm GeoChat GeoChat GeoChat TEOChatlas
GeoZero VHM-Instruct, fMoW, EuroSAT - RSVG, VHM-Instruct
wang2025geozero RSVQA-LR, WHU-RS19, DIOR-RSVG,
SkyEye968k, RSVQA-HR, RESISC45, VRSBench
VRSBench VRSBench NASC-TG2, AID VHM-Instruct
VHM-Instruct
Table 4: Data Pipeline of VLGFMs.
@@
liu2024remoteclip (the first 2 captions are generated according to the target location, and the remaining 3 captions are generated by
@@
zhang2024rs5m image checking and deduplication. Then, clean the dataset using the CLIP score and the RS image classifier. RS3:
@@
hu2023rsgpt type, and resolution; (2) object details such as quantity, color, shape, size, and position; (3) annotate the overall
@@
kuckreja2023geochat K-Means. Objects’ relative sizes are small, normal, or large. Objects’ relative positions are assessed using a 3x3
@@
mall2023remote ensuring a uniform location distribution avoiding duplicates, and filtering out indoor photos with a classifier.
@@
wang2024skyscript (OSM). Connect images with appropriate OSM semantics using a two-stage tag classification approach with
@@
zhan2024skyeyegpt Following the miniGPT-v2, create an phrase grounding dataset. Multi-task Conversation Instruction: mix the
@@
zhang2024earthgpt 1 visual grounding dataset into instructions using prompt templates.
@@
muhtar2024lhrsbot keys and balance semantic information. Using Vicuna-v1.5 to generate image captions by summarising the key
@@
pang2024h2rsvlm remote sensing images, including the type and resolution of the images, as well as information on the attributes,
@@
luo2024skysensegpt remote sensing objects.
GeoZero wang2025geozero GeoZero-Raw: Utilize the existing 13 RS vision-language datasets.
@@ 3.1 Data Pipeline
$\bullet$ Data collection from scratch. As shown in Table 3, three VLGFMs have built their training datasets from scratch rather than utilizing existing ones. RSGPT hu2023rsgpt prioritizes the quality of training samples over their quantity. Hence, it employs five experts to annotate images, producing high-quality captions manually. As a result, the model achieves satisfactory performance with fine-tuning on only 2,500 remote sensing image-text pairs. GRAFT mall2023remote ingeniously utilizes the geo-tag information of images to link ground-level image captions collected from social media with RSIs. This approach enables the acquisition of many remote sensing image-text pairs without the need for manual annotation of RSI captions. SkyScript wang2024skyscript utilizes a well-known open-source collaborative project, OpenStreetMap [^2] https://www.openstreetmap.org (OSM), to construct image captions. OSM allows users to contribute and edit geospatial data, such as road networks, trails, cafes, railway stations, etc. It provides a comprehensive and freely accessible world map and detailed user annotations. SkyScript filters these annotations and combines various tags to generate high-quality image captions. In LHRS-Bo muhtar2024lhrsbot, the LHRS-Align also acquires tags of geographic features from OSM and generates image captions based on these tags using language-only LLM, Vicuna-v1.5. More details can be found in Table 4.
@@
$\bullet$ Data enhancement using existing datasets. Most works choose to modify existing remote sensing datasets, which is significantly less labour-intensive compared to building datasets from scratch. Consequently, researchers can focus more on improving the model’s structures. RemoteCLIP liu2024remoteclip was the first to propose using text templates to convert the most common object detection annotations in the field of remote sensing interpretation into image captions, generating a large number of remote sensing image-text pairs suitable for training large models. Subsequently, EarthGPT zhang2024earthgpt also adopted the same approach. RS5M zhang2024rs5m, which currently contains the largest image-text pairs, includes 5 million RSIs. It employs the general multimodal model BLIP2 blip2 to generate image captions and then uses CLIP clip to select the top 5 highest-scoring captions. GeoChat kuckreja2023geochat constructs a remote sensing image-text pair dataset based on the segmentation dataset generated using SAM SAM technology. It first uses the target’s color, size, and location to build basic expressions and then inputs these expressions into a language-only LLM to generate question-answer pairs. SkyEyeGPT zhan2024skyeyegpt constructs a multi-task conversation instruction dataset by combining an object detection dataset with a VQA dataset built using the object detection data. This approach supports training for both VQA and visual grounding tasks. For more details, please refer to Table 4.
@@ 3.2 Architecture
VLGFMs can be briefly classified into three categories based on their input-output mechanisms. Contrastive VLGFMs liu2024remoteclip; zhang2024rs5m; mall2023remote; wang2024skyscript take both text and images as inputs and produce a similarity measure between them. This similarity is crucial for image-text retrieval and zero-shot scene classification applications. Conversational VLGFMs hu2023rsgpt; kuckreja2023geochat; zhan2024skyeyegpt; zhang2024earthgpt; muhtar2024lhrsbot; silva2024large; pang2024h2rsvlm also process text and images as inputs but output textual responses, leveraging the powerful language capabilities of LLMs to support tasks such as captioning and visual question answering. Additionally, dense prediction tasks can be supported by appending task-specific heads to the LLM. On the other hand, generative VLGFMs khanna2023diffusionsat; tang2024crsdiff are conditioned on either text or images, producing images as outputs. These models utilize conditional diffusion processes to generate controlled synthetic RSIs. This section will provide a detailed exposition of the model architectures for these three types of VLGFMs.
@@ 3.2.1 Contrastive VLGFMs
A typical contrastive VLGFM comprises two main components: an image encoder and a text encoder. A schematic of this architecture is depicted in Figure 2. Image encoder uses a convolutional neural network or a vision transformer to encode images into a feature vector. Parallel to the vision encoder, the text encoder (typically a BERT model) processes textual descriptions into another feature vector. The core idea behind contrastive VLGFM is to align the image and text embeddings in a shared multidimensional space. During training, contrastive VLGFM uses a contrastive learning objective, specifically a noise contrastive estimation technique variant. The model is trained to maximize the similarity between positive image-text pairs and minimize the similarity between negative pairs. The large-scale training allows contrastive VLGFM to understand various visual concepts described in natural language robustly. CLIP clip is one of the pioneering models for vision-language pre-training. It employs the InfoNCE loss to align text and images in the feature space. Benefiting from a dataset comprising 400 million image-text pairs, CLIP has achieved significant success in various multimodal downstream tasks. BLIP blip is the first multimodal model that simultaneously addresses both understanding and generation of image-text content, which is jointly pre-trained on three vision-language tasks: image-text contrastive learning, image-text matching, and image-conditioned language modelling.
In the domain of earth observation, some outstanding vision-language pre-training models have also emerged. RemoteCLIP liu2024remoteclip is the first true contrastive VLGFM. It utilizes existing remote sensing retrieval, detection, and segmentation datasets to construct image-text sample pairs, supporting model training. GeoRSCLIP is a vision-language pre-training model released alongside the RS5M dataset zhang2024rs5m. The authors did not provide extensive details about it in the paper. GRAFT mall2023remote is a significant attempt to address the issue of insufficient samples in remote sensing image-text data. It utilizes a vast collection of ground-level images from Flickr [^3] https://www.flickr.com as a bridge, pairing RSIs that correspond to the geographic coordinates of these ground images with the titles of the ground images. This approach allows for the construction of a large-scale dataset of remote sensing image-text pairs without the need for additional manual annotation. SkyScript wang2024skyscript utilizes the ground sample distance (GSD) of OSM data to annotate images appropriately. For instance, house numbers cannot be discerned from RSIs regardless of their resolution, and utility poles are not visible in RSIs with a GSD of 10. However, these contrastive VLGFMs are primarily data-centric efforts, with no significant modifications made to the model structures.
@@ 3.2.2 Conversational VLGFMs
Table 1 presents the visual encoders commonly used in current VLGFM implementations. Besides the standard CLIP image encoder clip, some studies have explored other variants. For instance, RSGPT and SkyEyeGPT employ the EVA-CLIP encoder sun2023evaclip, which is trained with improved training techniques. Additionally, EarthGPT utilizes two encoders simultaneously and introduces the ConvNext encoder convnext, a convolution-based architecture designed to exploit higher resolutions and multi-level features.
$\bullet$ Pre-trained LLM. Training an LLM from scratch entails prohibitive computational costs for most individuals, making it more effective and practical to begin with a pre-trained one. Through pre-training on vast amounts of data collected from various places, the LLM acquires rich world knowledge and demonstrates strong capabilities in generalization and reasoning. Table 1 summarizes commonly used and publicly available LLMs, all of which fall under the category of causal decoders gpt3. The LLAMA series llava and the Vicuna family vicuna represent some of the most notable open-source LLMs that have garnered considerable attention within the academic community. It is important to note that these LLMs were primarily pre-trained on English corpora, which limits their multilingual capabilities, such as for Chinese. Additionally, increasing the parameter size of LLMs also brings additional benefits, akin to increasing input resolution. Specifically, RSGPT hu2023rsgpt found that simply expanding the Vicuna from 7B to 13B parameters can significantly enhance performance across various benchmarks.
@@
For token-level connectors, the output of the encoder is transformed into tokens and concatenated with text tokens before being fed into the LLM. A common and feasible solution is to use a set of learnable query tokens to extract information in a query-based manner, initially implemented with BLIP-2 blip2 and subsequently adopted by various works on VLGFMs hu2023rsgpt. This query-based approach compresses visual tokens into a smaller number of representation vectors. In contrast, some methods employ multi-layer perception (MLP) to bridge the modal gap. For instance, the LLaVA series only use linear layers zhan2024skyeyegpt; zhang2024earthgpt; silva2024large or two-layer MLPs kuckreja2023geochat; pang2024h2rsvlm to project visual tokens and align feature dimensions with word embeddings.
@@
For feature-level connectors, it is required to incorporate additional modules to facilitate deep interaction and integration between textual features and visual features. For instance, Flamingo flamingo introduces additional cross-attention layers between the frozen transformer layers of the LLM, thereby utilizing external visual cues to enhance language features. Similarly, LHRS-Bot muhtar2024lhrsbot attempts to introduce a set of learnable queries for each level of image features. These queries aim to encapsulate the semantic information of each level of visual features through a series of stacked cross-attention and MLP layers.
@@
$\bullet$ Pre-training. Pre-training primarily aims to align different modalities and learn multimodal world knowledge in the initial training phase. The pre-training phase typically requires large-scale image-text pair data, such as caption data. For a given image, the model is trained to autoregressively predict the image’s caption using a standard cross-entropy loss. A common approach during pre-training is to keep pre-trained modules (such as the visual encoder and LLM) frozen and train the vision-language connector hu2023rsgpt. The idea is to align the visual and linguistic modalities without losing the knowledge acquired during pre-training. Some methods zhan2024skyeyegpt; zhang2024earthgpt; muhtar2024lhrsbot also unfreeze LLM to allow more trainable parameters for alignment. ShareGPT4V chen2023sharegpt4v found that having high-quality caption data during pre-training and unlocking visual encoding capabilities can promote better alignment. In the existing VLGFM landscape, only H2RSVLM pang2024h2rsvlm has attempted to unfreeze all modules simultaneously for pre-training.
@@
$\bullet$ Instruction-tuning. Instruction-tuning aims to teach models to understand user commands better and complete desired tasks. Through this approach, LLMs can be generalized to support unseen tasks, enhancing zero-shot performance. This simple yet effective idea has spurred the success of subsequent NLP works, such as InstructGPT instructgpt. Supervised fine-tuning methods typically require much training data to train task-specific models. Instruction-tuning methods reduce the reliance on large-scale data and can accomplish specific tasks through prompt engineering. Due to the flexible format of instructional data and the variability in templates across different tasks, collecting data samples is often more challenging and costly. Recent studies have shown that the quality of instructional fine-tuning samples is more critical than quantity. zeng2023matters discovered that high-quality instruction-tuning data can yield better performance, even if the quantity of samples is not the largest.